OpenStudy (anonymous):

Use the distributive property to simplify the expression below:4x-8/4

OpenStudy (swag):

4x-2

OpenStudy (anonymous):

\[\left( 4x -8 \right)\div4 = \left( 4x \div4 \right)-\left( 8\div4 \right) = x-2\]